mnjones1:

A diver dives 17 and two-thirds yards and then comes back to the surface. Then, she dives 9 and StartFraction 3 over 8 EndFraction yards deeper than her first dive. What is her depth, in yards, in relation to the surface of the water? Negative 27 and StartFraction 7 over 24 EndFraction Negative 27 and StartFraction 1 over 24 EndFraction Negative 8 and StartFraction 7 over 24 EndFraction Negative 8 and StartFraction 1 over 24 EndFraction

Mercury:

The first dive is \[-17\frac{ 2 }{ 3 }\](negative because it's underneath sea level) she dove another \[9\frac{ 3 }{ 8 }\] which means her depth is now \[-17\frac{ 2 }{ 3 }-9\frac{ 3 }{ 8 }\] now, ignore the negative signs for now, and just focus on converting the mixed numbers to improper fractions. multiply the denominator * the whole number part, then add the numerator, then put the result over the original denominator (easier to see through an example) 

Mercury:

repeating this logic with 9 and 3/8, we get 75/8 now we have 53/3 and 75/8 to convert these both to the same denominator, we find the LCM of 3 and 8 (which is 24) so we multiply 53/3 by 8/8 to get 424/24 and 75/8 by 3/3 to get 225/24 then adding in the original signs we have -424/24 - 225/24 = -649/24 24 goes into 649, 27 times with 1 leftover ----> -649/24 = -27 and 1/24